OpenStudy (praxer):

Probability: A card is drawn from a pack. The card is replaced and the pack is reshuffled. If this is done six times , the probability that 2 hearts, 2 diamonds and 2 black cards are drawn is. 90/2^(10) I am unable to get the result. Please help !!!

OpenStudy (praxer):

@ganeshie8

OpenStudy (kittiwitti1):

So the result is supposed* to be this?\[\left(\frac{90}{2}\right)^{10}\]

OpenStudy (praxer):

90/(2^(10))

OpenStudy (welshfella):

probability of a heart being drawn in 1 draw = 1/4 for a diamond = 1/4 for a black card = 1/2 probability of drawing HHDDBB = 1/4 * 1/4 *1/4*1/4*1/2*1/2= 1 / 1024

OpenStudy (welshfella):

which is 1 / 2^10

OpenStudy (kittiwitti1):

Sorry, my internet lagged out. So this?\[\left(\frac{90}{2^{10}}\right)\]

OpenStudy (welshfella):

- but this result is just for the given order The order of drawing could be HHBDBD for example

OpenStudy (praxer):

yes it is @kittiwitti1

OpenStudy (kittiwitti1):

Okay. Well, I would try to help but @welshfella seems to be helping you already, so I'll just go... ; w;

OpenStudy (welshfella):

so we need to work out how many different arrangements of the cards are possible - and hope its 90!

ganeshie8 (ganeshie8):

90! is an astronomical number =P

OpenStudy (welshfella):

I think its 6! / 2^3

OpenStudy (welshfella):

6! / 2! 2! 2!

OpenStudy (welshfella):

- if my memory of this stuff serves me well..

OpenStudy (praxer):

6C2*4C2*2C2*1/4 * 1/4 *1/4*1/4*1/2*1/2=90/(2^10) will it be like this. Cause for 2 hearts to obtain of 6 cards is 6C2 and from the rest 4 we need 2 diamonds and 2 black cards obtained at the last.

OpenStudy (welshfella):

= 720/8 = 90

OpenStudy (praxer):

6C2 * 4C2 * 2C2 =90

ganeshie8 (ganeshie8):

both will work, bt welshfella's method looks a lot simpler

OpenStudy (welshfella):

yes - your method is valid @praxer

ganeshie8 (ganeshie8):

`HHDDBB` possible permutations = 6!/(2!2!2!)

OpenStudy (welshfella):

lol when I said 90! I didn't mean factorial 90 IIt was just an exclamation mark - I didnt see that at the time!!!

ganeshie8 (ganeshie8):

i was jk haha

OpenStudy (praxer):

thank you @welshfella and @ganeshie8 :)

OpenStudy (welshfella):

oh ok! - I like using exclamation marks...

OpenStudy (welshfella):

yw